# BMRetriever: Tuning Large Language Models as Better Biomedical Text Retrievers

### v1

@@ 1 Introduction
In the field of biomedicine, the ability to effectively retrieve knowledge from external corpora is crucial for large language models (LLMs) to excel at biomedical NLP tasks (Lewis et al., 2020; Zhang et al., 2024b; Asai et al., 2024). By tapping into up-to-date domain knowledge, retrieval-augmented LLMs have demonstrated promising results in various biomedical downstream applications, including knowledge discovery (Frisoni et al., 2022), question answering (Wang et al., 2023, 2024a), and clinical decision-making (Naik et al., 2022; Shi et al., 2023; Xu et al., 2024).
@@
Several works have designed specialized retrieval models for biomedical domains  Mohan et al. (2017); Liu et al. (2021); Jin et al. (2023); Luo et al. (2022a); Singh et al. (2023); Zhang et al. (2023). However, these models are typically built upon BERT-series models, which have limited representative power. Besides, they often rely on proprietary data (e.g., private search logs or patient records), making it challenging to scale them up to accommodate larger models effectively due to privacy concerns. While recent studies in the general domain have improved neural retrieval models via scaling up model size Ni et al. (2022); Muennighoff (2022); Wang et al. (2024b) and training data (Izacard et al., 2022; Wang et al., 2022b; Lin et al., 2023), adapting such models to the biomedical domain may lead to suboptimal performance due to the distribution shift issue (Thakur et al., 2021). Developing large-scale retrieval models dedicated to the biomedical domain without requiring massive proprietary datasets remains crucial yet challenging.
@@
We conduct extensive experiments across five tasks on eleven biomedical datasets to demonstrate the strong performance of BMRetriever. As shown in Figure 1, BMRetriever outperforms existing dense retrievers with orders of magnitude more parameters: with 410M parameters, it surpasses the performance of GTR-4.8B (Ni et al., 2022) and SGPT-2.7B (Muennighoff, 2022), which have 7×7$\times$7 × more parameters. At the 7B scale, BMRetriever outperforms the recently proposed E5-Mistral (Wang et al., 2024b), which uses extra-large batch-size and nonpublic data mixture. In addition, BMRetriever presents a lightweight yet high-performing domain adaptation solution, with its 1B variant achieving more than 98% performance of E5-Mistral using only 14.3% of parameters. Our contribution can be summarized as follows:
@@ 2 Related Work
Earlier research explores various approaches for learning representations suitable for text retrieval (Deerwester et al., 1990; Huang et al., 2013). More recently, several studies introduce dual-encoder architectures based on BERT for dense retrieval Karpukhin et al. (2020); Xiong et al. (2021); Qu et al. (2021); Gao and Callan (2022); Izacard et al. (2022); Xiao et al. (2022). With the advent of LLMs with billions of parameters, several studies attempt to scale up model size (Ni et al., 2022; Neelakantan et al., 2022; Muennighoff, 2022; Springer et al., 2024), often fine-tuned on multi-task instruction data (Asai et al., 2023; Su et al., 2023; Wang et al., 2024b; Muennighoff et al., 2024). However, the benefit of scaling up is more pronounced for general domain datasets where massive annotated data are available.
To design effective retrievers for specialized domains, several works propose continuously pre-train the retrieval model on domain-specific corpora (Yu et al., 2022; Zhang et al., 2023) or fine-tuning the model on proprietary search datasets (Mohan et al., 2017; Jin et al., 2023). On the other hand, synthetic data has also been used to improve the generalization ability of dense retrieval model (Ma et al., 2021; Wang et al., 2022a; Jiang et al., 2023; Wang et al., 2024b; Zhang et al., 2024a). Despite these advancements, how to combine public, open data to formulate a dataset curation recipe for adapting LLMs as high-performing biomedical retrievers remains unresolved. Our method efficiently integrates diverse supervision signals for biomedical retrieval model training, which achieves better performance than baselines trained with more data.
@@ 3.2 Unsupervised Contrastive Pre-training
Contrastive pre-training improves the quality of representations by better aligning similar text sequences while ensuring the uniformity of unrelated text sequences, which helps adapt the retrieval model to biomedical domains (Gururangan et al., 2020; Yu et al., 2022; Luo et al., 2022b).
@@ 3.3 Supervised Instruction Fine-tuning
Beyond relevance signals, task generalization is also crucial for building a general retriever, as user intent and input formats vary while public data captures only a fraction of tasks. To address this, we perform task-level augmentation, which involves prompting GPT-4 (gpt-4-turbo-1106) to conceptualize a diverse list of potential scenarios for biomedical retrieval tasks (Wang et al., 2024b). Subsequently, we prompt GPT-4 again to generate examples for each scenario, including a query, a relevant (positive) passage, and a challenging irrelevant (hard negative) passage. This approach allows us to enhance the diversity of instructions.
Hard Negative Mining and Data Filter. In both labeled instruction fine-tuning datasets and data-label synthetic datasets, positive pairs are available, while negative examples are missing. To obtain the negatives, we randomly select 1 passage from the top 100 passages retrieved by E5-base (Wang et al., 2022b) when using the given query to search the entire corpus of the corresponding dataset. As the generated synthetic data can be noisy, consistency filtering is adopted to filter low-quality pairs (Alberti et al., 2019; Dai et al., 2023), where for each synthetic (query $q$, passage $p$) pair, we use the E5-base to predict the most relevant passages for ${q}$. We only retain $q$ when $p$ occurs among the top three retrieved passages.
@@ 4.1 Experimental Setups
Baselines. We compare to sparse retrieval models BM25 (Robertson et al., 2009) and open-source dense retrieval models with varying model sizes: Contriever (Izacard et al., 2022), Dragon (Lin et al., 2023), SciMult (Zhang et al., 2023), SPECTER 2.0 (Singh et al., 2023), COCO-DR (Yu et al., 2022), QExt (Meng et al., 2022), SGPT (Muennighoff, 2022), MedCPT (Jin et al., 2023), GTR (Ni et al., 2022), InstructOR (Su et al., 2023), E5-Large-v2 (Wang et al., 2022b), BGE-Large (Chen et al., 2024), LLaRA (Li et al., 2023a), RepLLaMA (Ma et al., 2023), Grit-LM (Muennighoff et al., 2024), LLM2Vec (BehnamGhader et al., 2024), E5-Mistral (Wang et al., 2024b), and CPT-text (Neelakantan et al., 2022). The details of baselines and parameter sizes are in Appendix C.
Evaluation. We follow existing works (Muennighoff et al., 2023; Thakur et al., 2021) and use nDCG@10 to measure standard IR performance and Spearman correlation for STS based on cosine similarity. To evaluate the retrieval performance of QA, we report Recall@{5,20} and nDCG@20. For entity linking, we report mean reciprocal rank (MRR)@5 and Recall@{1,5}. For paper recommendation, we follow Singh et al. (2023) and report mean average precision (MAP) and nDCG. The implementation details can be found in Appendix D.
@@ 4.3 Results on Retrieval-Oriented Biomedical Applications
4.4 Studies on Contrastive Pre-training
@@
To highlight the effectiveness of our contrastive pre-training approach, we evaluate the performance of unsupervised dense retrieval models that only use unlabeled corpora for pre-training. As shown in Table 4, our model outperforms existing unsupervised baselines and even surpasses many fully supervised models reported in Table 2. The strong unsupervised results have important implications for real-world biomedical applications, where curating large labeled datasets is often prohibitively expensive and time-consuming. Our approach presents an attractive alternative, enabling the development of high-quality retrieval models in a data-efficient manner using only the unlabeled corpus.
@@ 4.5 Studies on Instruction Fine-tuning
Figure 3 illustrates the impact of different fine-tuning data sources on model performance across various datasets [^2] Removing biomedical data retains the synthetic data.. Among all the utilized data types, synthetic data contributes the most significant performance gain, which can be attributed to its larger volume compared to biomedical data and its coverage of a more diverse range of task types. It is particularly beneficial for NFCorpus, SciFact, and Trec-COVID, as these datasets follow the standard IR format of short queries and long passages, aligning with the format of the synthetic data. Furthermore, synthetic data proves advantageous for the iCliniq dataset, as it potentially includes various retrieval scenarios, such as dialog data. General domain fine-tuning data, consisting of short queries and long passages, generally enhances relevance estimation capabilities, benefiting standard biomedical IR tasks like Trec-COVID. However, it may slightly diminish performance on datasets like iCliniq, which contain conversational patient-provider data that deviates from traditional IR formats, potentially introducing distribution shifts. Biomedical fine-tuning data, on the other hand, demonstrates broad usefulness, particularly for BIOSSES and iCliniq tasks, likely due to its inclusion of sentence similarity and dialog data aligning with downstream task formats. Public fine-tuning data from E5-Mistral (Wang et al., 2024b) and MEDI, used by InstructOR (Su et al., 2023), achieves relatively poor performance, possibly due to their focus on the general domain with limited biomedical knowledge. Additionally, the lack of publicly available synthetic generated data from E5-Mistral (Wang et al., 2024b) may contribute to the suboptimal performance observed.
@@ 5 Conclusion
Cost Estimation. Generating synthetic data using GPT models incurs additional costs. In our work, the total API cost of BMRetriever is less than \$500444As of April 2024., which remains affordable within an academic budget. This cost is significantly lower than recent works (Wang et al., 2024b), which have an estimated cost of more than \$6000.
Other Training Objective and Backbone Models. The primary focus of this work is on building an effective dataset creation recipe to scale up retrieval models for the biomedical domain. However, other biomedical LLMs can also be used as the backbones (Luo et al., 2022b; Bolton et al., 2024; Chen et al., 2023). In our early attempts, we did not observe significant performance gains with some of these backbones (e.g., BioGPT-Large, BioMedLM-2.7B, Meditron-7B), potentially due to the loose connection between the objective of sequence contrastive learning and causal language modeling. Besides, since we utilize standard contrastive learning objectives with hard negative mining without proposing new learning objectives, we do not claim this as our core contribution. Exploring ways to leverage domain-specific LLMs with improved training techniques to enhance performance is an interesting topic for future research.
@@
Misinformation. One specific issue for LLM-generated biomedical text is the potential for misinformation and hallucination (Pal et al., 2023). It is important to note that for the generated queries, the majority are short sentences or phrases without presenting any scientific facts. Regarding the generated (query, passage) pairs, to ensure that our generated synthetic text does not introduce misinformation or hallucination, we randomly selected 200 examples and asked medical students to evaluate the factuality of the generated text. The evaluation results did not reveal any instances of misinformation or hallucination in the selected examples.
Data Contamination. A potential issue is test set contamination (Sainz et al., 2023), where some test examples overlap with the training data. This can be especially problematic for text generated by LLMs, as they are often pre-trained on massive corpora spanning various domains. To address this concern, we follow Wang et al. (2024b) to conduct a string match-based analysis between the test set and our training set, where we do not observe any overlap between the train and test queries. While some of the corpora (e.g., PubMed) are also utlized in the test tasks, this is a standard practice even in zero-shot or few-shot evaluation of retrieval models (Ma et al., 2021; Wang et al., 2022a; Yu et al., 2022; Dai et al., 2023), and it is not considered as contamination.
@@ B.1 Pre-training Corpus
Dataset Size Line
PubMed Xiong et al. (2024) 8M https://huggingface.co/datasets/MedRAG/pubmed
arXiv, MedRxiv, BioRxiv Muennighoff et al. (2023) 577K https://huggingface.co/datasets/mteb/raw_arxiv
Meadow Wang et al. (2020) 460k https://huggingface.co/datasets/medalpaca/medical_meadow_cord19
Textbooks Jin et al. (2021) 50K https://huggingface.co/datasets/MedRAG/textbooks
StatPearls Xiong et al. (2024) 54K https://huggingface.co/datasets/MedRAG/statpearls
LitCovid Chen et al. (2021) 70K https://huggingface.co/datasets/KushT/LitCovid_BioCreative
S2ORC Lo et al. (2020) 600K https://github.com/allenai/s2orc
MS Marco Bajaj et al. (2016) 1.2M https://huggingface.co/datasets/Tevatron/msmarco-passage-corpus
@@ C.1 Baselines for Retrieval Tasks in Main Experiments
  • Contriever (Izacard et al., 2022) is a dense retrieval model (110M) pre-trained via contrastive learning on documents sampled from Wikipedia and CC-Net (Wenzek et al., 2020) corpora.
@@
  • Grit-LM (Muennighoff et al., 2024) is a generative dense retrieval model (7B) that incorporates generative representational instruction tuning, enabling the model to handle both generative and embedding tasks by distinguishing between them through instructions.
  • E5-Mistral (Wang et al., 2024b) is an enhanced version of the E5 (Wang et al., 2022b) that incorporates synthetic data generated by LLMs for a diverse range of text embedding tasks.
@@ C.2 Baselines for Retrieval-Oriented Downstream Applications
In experiments for retrieval-oriented downstream applications, we only compare BMRetriever to the strongest, most relevant, and fair baselines, including: (1) Base Size (<1B): Dragon (Lin et al., 2023), MedCPT (Jin et al., 2023), and E5-Large-v2 (Wang et al., 2022b); (2) Large Size (1B-5B): InstructOR (Su et al., 2023) and SGPT-2.7B (Muennighoff, 2022); and (3) XL Size (>5B): E5-Mistral (Wang et al., 2024b).

### v2

@@ 1 Introduction
In the field of biomedicine, the ability to effectively retrieve knowledge from external corpora is crucial for large language models (LLMs) to excel at biomedical NLP tasks (Lewis et al., 2020). By tapping into up-to-date domain knowledge, retrieval-augmented LLMs have demonstrated promising results in various biomedical downstream applications, including knowledge discovery (Frisoni et al., 2022), question answering (Zhang et al., 2024b; Yu et al., 2024), and clinical decision-making (Naik et al., 2022; Shi et al., 2023; Xu et al., 2024).
@@
Several works have designed specialized retrieval models for biomedical domains  Mohan et al. (2017); Liu et al. (2021); Jin et al. (2023); Luo et al. (2022); Singh et al. (2023); Zhang et al. (2023). However, these models are typically built upon BERT-series models, which have limited representative power. Besides, they often rely on proprietary data (e.g., private search logs or patient records), making it challenging to scale them up to accommodate larger models effectively due to privacy concerns. While recent studies in the general domain have improved neural retrieval models via scaling up model size Ni et al. (2022); Wang et al. (2024) and training data (Izacard et al., 2022; Wang et al., 2022b; Lin et al., 2023), adapting such models to the biomedical domain may lead to suboptimal performance due to the distribution shift issue (Thakur et al., 2021). Developing large-scale retrieval models dedicated to the biomedical domain without requiring massive proprietary datasets remains crucial yet challenging.
@@
We conduct extensive experiments across five tasks on eleven biomedical datasets to demonstrate the strong performance of BMRetriever. As shown in Figure 1, BMRetriever outperforms existing dense retrievers with orders of magnitude more parameters: with 410M parameters, it surpasses the performance of GTR-4.8B (Ni et al., 2022) and SGPT-2.7B (Muennighoff, 2022), which have 7×7$\times$7 × more parameters. At the 7B scale, BMRetriever outperforms the recently proposed E5-Mistral (Wang et al., 2024), which uses extra-large batch-size and nonpublic data mixture. In addition, BMRetriever presents a lightweight yet high-performing domain adaptation solution, with its 1B variant achieving more than 98% performance of E5-Mistral using only 14.3% of parameters. Our contribution can be summarized as follows:
@@ 2 Related Work
Earlier research explores various approaches for learning representations suitable for text retrieval (Deerwester et al., 1990; Huang et al., 2013). More recently, several studies introduce dual-encoder architectures based on BERT for dense retrieval Karpukhin et al. (2020); Xiong et al. (2021); Gao and Callan (2022); Izacard et al. (2022). With the advent of LLMs with billions of parameters, several studies attempt to scale up model size (Ni et al., 2022; Neelakantan et al., 2022), often fine-tuned on multi-task instruction data (Asai et al., 2023; Su et al., 2023; Wang et al., 2024). However, the benefit of scaling up is more pronounced for general domain datasets where massive annotated data are available.
To design effective retrievers for specialized domains, several works propose continuously pre-train the retrieval model on domain-specific corpora (Yu et al., 2022; Zhang et al., 2023) or fine-tuning the model on proprietary search datasets (Mohan et al., 2017; Jin et al., 2023). On the other hand, synthetic data has also been used to improve the generalization ability of dense retrieval model (Wang et al., 2022a; Jiang et al., 2023; Wang et al., 2024; Zhang et al., 2024a). Despite these advancements, how to combine public, open data to formulate a dataset curation recipe for adapting LLMs as high-performing biomedical retrievers remains unresolved. Our method efficiently integrates diverse supervision signals for biomedical retrieval model training, which achieves better performance than baselines trained with more data.
@@ 3.2 Unsupervised Contrastive Pre-training
Contrastive pre-training improves the quality of representations by better aligning similar text sequences while ensuring the uniformity of unrelated text sequences, which helps adapt the retrieval model to biomedical domains (Gururangan et al., 2020; Yu et al., 2022).
@@ 3.3 Supervised Instruction Fine-tuning
Beyond relevance signals, task generalization is also crucial for building a general retriever, as user intent and input formats vary while public data captures only a fraction of tasks. To address this, we perform task-level augmentation, which involves prompting GPT-4 (gpt-4-turbo-1106) to conceptualize a diverse list of potential scenarios for biomedical retrieval tasks (Wang et al., 2024). Subsequently, we prompt GPT-4 again to generate examples for each scenario, including a query, a relevant (positive) passage, and a challenging irrelevant (hard negative) passage. This approach allows us to enhance the diversity of instructions.
Hard Negative Mining and Data Filter. In both labeled instruction fine-tuning datasets and data-label synthetic datasets, positive pairs are available, while negative examples are missing. To obtain the negatives, we randomly select 1 passage from the top 100 passages retrieved by E5-base (Wang et al., 2022b) when using the given query to search the entire corpus of the corresponding dataset. As the generated synthetic data can be noisy, consistency filtering is adopted to filter low-quality pairs (Alberti et al., 2019), where for each synthetic (query $q$, passage $p$) pair, we use the E5-base to predict the most relevant passages for ${q}$. We only retain $q$ when $p$ occurs among the top three retrieved passages.
@@ 4.1 Experimental Setups
Baselines. We compare to sparse retrieval models BM25 (Robertson et al., 2009) and open-source dense retrieval models with varying model sizes: Contriever (Izacard et al., 2022), Dragon (Lin et al., 2023), SciMult (Zhang et al., 2023), SPECTER 2.0 (Singh et al., 2023), COCO-DR (Yu et al., 2022), QExt (Meng et al., 2022), SGPT (Muennighoff, 2022), MedCPT (Jin et al., 2023), GTR (Ni et al., 2022), InstructOR (Su et al., 2023), E5-Large-v2 (Wang et al., 2022b), BGE-Large (Chen et al., 2024), LLaRA (Li et al., 2023a), RepLLaMA (Ma et al., 2023), LLM2Vec (BehnamGhader et al., 2024), E5-Mistral (Wang et al., 2024), and CPT-text (Neelakantan et al., 2022). The details of baselines and parameter sizes are in Appendix C.
Evaluation. We use nDCG@10 to measure standard IR performance and Spearman correlation for STS based on cosine similarity. To evaluate the retrieval performance of QA, we report Recall@{5,20} and nDCG@20. For entity linking, we report mean reciprocal rank (MRR)@5 and Recall@{1,5}. For paper recommendation, we follow Singh et al. (2023) and report mean average precision (MAP) and nDCG. The implementation details can be found in Appendix D.
@@ 4.3 Results on Retrieval-Oriented Biomedical Applications
4.4 Unsupervised Retrieval Performance
@@
To highlight the effectiveness of our contrastive pre-training approach, we evaluate the performance of unsupervised dense retrieval models that only use unlabeled corpora for pre-training and synthetic data for finetuning. As shown in Table 4, our model outperforms existing unsupervised baselines and even surpasses many fully supervised models reported in Table 2. The strong unsupervised results have important implications for real-world biomedical applications, where curating large labeled datasets is often prohibitively expensive and time-consuming. Our approach presents an attractive alternative, enabling the development of high-quality retrieval models in a data-efficient manner.
@@ 4.5 Studies on Instruction Fine-tuning
Figure 3 illustrates the impact of different fine-tuning data sources on model performance across various datasets [^2] Removing biomedical data retains the synthetic data.. Among all the utilized data types, synthetic data contributes the most significant performance gain, which can be attributed to its larger volume compared to biomedical data and its coverage of a more diverse range of task types. It is particularly beneficial for NFCorpus, SciFact, and Trec-COVID, as these datasets follow the standard IR format of short queries and long passages, aligning with the format of the synthetic data. Furthermore, synthetic data proves advantageous for the iCliniq dataset, as it potentially includes various retrieval scenarios, such as dialog data. General domain fine-tuning data, consisting of short queries and long passages, generally enhances relevance estimation capabilities, benefiting standard biomedical IR tasks like Trec-COVID. However, it may slightly diminish performance on datasets like iCliniq, which contain conversational patient-provider data that deviates from traditional IR formats, potentially introducing distribution shifts. Biomedical fine-tuning data, on the other hand, demonstrates broad usefulness, particularly for BIOSSES and iCliniq tasks, likely due to its inclusion of sentence similarity and dialog data aligning with downstream task formats. Public fine-tuning data from E5-Mistral (Wang et al., 2024) and MEDI, used by InstructOR (Su et al., 2023), achieves relatively poor performance, possibly due to their focus on the general domain with limited biomedical knowledge. Additionally, the lack of publicly available synthetic generated data from E5-Mistral (Wang et al., 2024) may contribute to the suboptimal performance observed.
@@ 5 Conclusion
Cost Estimation. Generating synthetic data using GPT models incurs additional costs. In our work, the total API cost of BMRetriever is less than \$500444As of April 2024., which remains affordable within an academic budget. This cost is significantly lower than recent works (Wang et al., 2024), which have an estimated cost of more than \$6000.
@@
Misinformation. One specific issue for LLM-generated biomedical text is the potential for misinformation and hallucination (Pal et al., 2023). It is important to note that for the generated queries, the majority are short sentences or phrases without presenting any scientific facts. We randomly selected 200 examples and asked medical students to evaluate the factuality of the generated text. The evaluation results did not reveal any misinformation or hallucination in the selected examples.
Data Contamination. A potential issue is test set contamination (Sainz et al., 2023), where some test examples overlap with the training data. This can be especially problematic for text generated by LLMs, as they are often pre-trained on massive corpora spanning various domains. To address this concern, we follow Wang et al. (2024) to conduct a string match-based analysis between the test set and our training set, where we do not observe any overlap between the train and test queries. While some of the corpora (e.g., PubMed) are also utilized in the test tasks, this is a standard practice even in zero-shot or few-shot evaluation of retrieval models (Wang et al., 2022a; Yu et al., 2022), and it is not considered as contamination.
Acknowledgement
We thank the anonymous reviewers and area chairs for valuable feedbacks. This research was also partially supported by the National Science Foundation under Award Number 2319449 and Award Number 2312502, the National Institute Of Diabetes And Digestive And Kidney Diseases of the National Institutes of Health under Award Number K25DK135913, the Emory Global Diabetes Center of the Woodruff Sciences Center, Emory University. JH was supported by NSF grants IIS-1838200 and IIS-2145411. YY and CZ was supported in part by NSF IIS-2008334 and CAREER IIS-2144338.
@@ C.1 Baselines for Retrieval Tasks in Main Experiments
  • Contriever (Izacard et al., 2022) is a dense retrieval model (110M) pre-trained via contrastive learning on documents sampled from Wikipedia and CC-Net corpora.
@@
  • E5-Mistral (Wang et al., 2024) is an enhanced version of the E5 (Wang et al., 2022b) that incorporates synthetic data generated by LLMs for a diverse range of text embedding tasks.
@@ C.2 Baselines for Retrieval-Oriented Downstream Applications
In experiments for retrieval-oriented downstream applications, we only compare BMRetriever to the strongest, most relevant, and fair baselines, including: (1) Base Size (<1B): Dragon (Lin et al., 2023), MedCPT (Jin et al., 2023), and E5-Large-v2 (Wang et al., 2022b); (2) Large Size (1B-5B): InstructOR (Su et al., 2023) and SGPT-2.7B (Muennighoff, 2022); and (3) XL Size (>5B): E5-Mistral (Wang et al., 2024).